CHRISTOPHER B. D’SOUZA

Department of Economics (613) 533-2276 (work/voicemail) Queen’s University (613) 545-1026 (home) Kingston, Ontario (613) 533-6668 (fax) Canada, K7L 3N6 e-mail: dsouza@qed.econ.queensu.ca

EDUCATION

Candidate for Ph.D. in Economics, Queen’s University, 1993 to Present • Fields of specialization: international economics and industrial organization • Other teaching fields: macroeconomics and monetary economics

• Thesis title: A Monetary Approach to the Exchange Rate

• Expected completion date: April 1997

Masters of Arts in Economics, Queen’s University, October 1992

• Field of specialization: industrial organization

• M.A. Essay title: Migration Flows and Regional Wage Adjustment

Bachelor of Arts (Honours), University of Western Ontario, May 1991

AWARDS

• Queen’s Graduate Fellowship: 1991-1994, 1995-1996

• Ontario Graduate Scholarship: 1994-1995

• S.F. Glass Gold Medal in Economics, U.W.O., 1991

• Dean’s Honour List: 1988-1991

• In-Course Scholarships: 1988-1991

EMPLOYMENT

Queen’s University, Department of Economics

• January-April 1997 Instructor: international trade

• 1995-1996 Instructor: introduction to UNIX

• 1993-1996 Teaching Assistant: graduate macroeconomics

• 1992-1993 Research Assistant: exchange rate management

• 1991-1992 Teaching Assistant: Canadian economic policy

Bank of Canada, Research Department

• 1990 Research Assistant: industrial regional analysis

University of Western Ontario, Faculty of Social Sciences

• 1989-1991 Computer Advisor PAPERS

• “Learning in Foreign Exchange Markets” 1995 mimeo

• “A New Monetary Model of the Exchange Rate” 1996 mimeo

• “Forecasting the Exchange Rate with a Monetary Model” 1996 mimeo

CONFERENCE PRESENTATIONS

• 1996 Canadian Economic Association Meetings:

“A New Monetary Model of the Exchange Rate”

PROFESSIONAL ACTIVITIES

• Conference Discussant: 1996 Canadian Economic Association Meetings • Referee: Journal of International Economics

CITIZENSHIP

• Canadian

REFERENCES

• Professor Gregor W. Smith (613) 533-6659, smithgw@qed.econ.queensu.ca

• Professor Beverly Lapham (613) 545-2297, laphamb@qed.econ.queensu.ca

• Professor Charles Beach (613) 545-2278, beachc@qucdn.queensu.ca

Department of Economics, Queen’s University Kingston, Ontario, Canada, K7L 3N6

PH.D. DISSERTATION ABSTRACT

The first essay of my thesis examines prices, returns and trading volumes in asset markets when agents are assumed to have expectations that are not fully rational. A dynamic model of speculative trading is presented in which agents have private, but imperfect, information about the fundamentals of the economy. In the bounded rational environment, agents continue to learn about the true probability distribution of the model. The analysis focuses on empirical regularities of the foreign exchange market. In contrast to the notrade implication of the rational expectations equilibrium, positive trading persists in the learning environment of the model. Results suggest that rational expectations and bounded rational environments have very similar time-series properties. Both display fat tails in returns, but neither environment provides evidence of non-stationary prices or volatility clustering in returns. Lastly, while rational expectations prices follow a similar pattern as the underlying fundamentals of the economy, bounded rational prices do not.

The second essay presents a monetary model of the exchange rate that, unlike most models in this class, provides an explicit role for trading in foreign exchange markets. Utility maximizing agents choose an amount of wealth to be invested in foreign assets that is determined partly by private information about domestic monetary policy and partly by risk and return factors. Differences in information about monetary policies among agents generate trade and realistic exchange rate dynamics. Implications of the model include nominal exchange rates with unit roots, volatility clustering, and fat-tails in exchange rate returns. Further, a calibration exercise illustrates that there is a close correspondence between the time series properties of our model and the historical sample.

The third essay reconsiders the previous model and evaluates the results of several published studies concerned with the out-of-sample forecasting accuracy of the structural monetary model of the exchange rate. We demonstrate, using artificial data on exchange rates, that poor forecasting results may arise from the improper modelling of expectations formation and non-linearities in the true data generating process. We consider the out-of-sample predictive performance of four models. Our results are qualitatively similar to published studies that employ actual time-series data on nominal exchange rates. Specifically, we find that a long-horizon model performs better that the random walk model, which in turn performs better that the a partial adjustment money demand model. The static monetary model is the poorest performer in out-of-sample forecasting performance.